# Nodal superconducting gap structure and topological surface states of UTe2

### v1

@@ I Introduction
The heavy-fermion compound UTe2 is currently a prime candidate for featuring an intrinsic spin-triplet superconducting phase in its ground state [3, 4, 5]. This is most evident from large critical magnetic fields for the destruction of superconductivity, well beyond the Pauli limiting field, and the weak temperature dependence of the Knight shift upon entering the superconducting phase [3, 6, 7, 8, 9, 10]. UTe2 is a boby-centered orthorhombic material with D2h point group symmetry, allowing four symmetry-distinct spin-triplet order parameters in the case of strong spin-orbit coupling (SOC): Au, B1u, B2u, and B3u. Time-reversal symmetry can be broken below $T_{c}$ in the case of accidental near-degeneracies leading to superpositions of these states, as for example B1u + iB2u. At present, the experimental status of the nature of superconductivity in UTe2 remains controversial, with evidence in favor of all four allowed irreducible representations (irreps) and some also advocating for complex time-reversal symmetry-breaking non-unitary superpositions [11]. Thus, further experiments and theoretical modeling are needed to pinpoint the nature of the spin-triplet ordering of this material.
Several recent developments driven by improved sample quality appear to pave the way for significant progress in the understanding of superconductivity in UTe2 [12]. The superconducting state is not reached by multiple thermodynamic transitions and time-reversal symmetry remains preserved in the ground state condensate [13, 14]. It appears that the earlier signatures of time-reversal symmetry breaking below $T_{c}$ originated from extrinsic sources [15, 16, 17, 18]. Thus, the superconducting state most likely resides within a single of the above-mentioned irreps [19]. Another important development is in the description of the normal state low-energy electronic states and their associated Fermi surface. Quantum oscillations are consistent with corrugated cylindrical Fermi sheets [20, 21]. Evidence for a closed Fermi pocket [22, 23, 24] is relatively weak, and hence we assume in this work the absence of closed Fermi pockets, which is important for the determination of the leading pairing instability and its associated topological properties [25, 26, 27, 28, 29]. Given the extensive evidence for nodal quasiparticles from e.g. thermal conductivity [30, 31], specific heat [32, 33] and penetration depth measurements [11], the leading candidates for the superconducting order parameter appear to be B2u and B3u.
@@ II Methodology
In this section, we discuss first the applied normal state band structure which is based on a slightly modified version of the density functional theory (DFT) calculations and subsequent four-band tight-binding model presented in Ref. [19]. Second, we describe the derivation of the symmetry-allowed superconducting terms for both uranium (U) and tellurium (Te) sublattice sites. Finally, we outline the method used for extracting the surface Majorana bound states inside the superconducting gap.
@@ II.1 Model for the normal state
The starting point is a tight-binding model with few neighbor hoppings as used in Ref. [19] to describe the electronic structure and Fermi surface, consistent with recent conclusions from quantum oscillation measurements [20, 21]. The Hamiltonian of the U degrees of freedom can be expanded in Pauli matrices $\rho_{i}$ as
@@
$-\rho_{1}\;$ $[\Delta_{\mathrm{U}}+2t_{U}^{\prime}\cos k_{x}+2t_{\mathrm{ch,U}}^{\prime}\cos k_{y}$
@@
$\equiv\phantom{-}\rho_{0}\;$ $f_{\mathrm{U},0}(\mathbf{k})+\rho_{1}\;f_{\mathrm{U},1}(\mathbf{k})+\rho_{2}\;f_{\mathrm{U},2}(\mathbf{k}),$ (1)
@@
$\equiv\phantom{-}\rho_{0}\;$ $f_{\mathrm{Te},0}(\mathbf{k})+\rho_{1}\;f_{\mathrm{Te},1}(\mathbf{k})+\rho_{2}\;f_{\mathrm{Te},2}(\mathbf{k}).$ (2)
@@
$H_{N}(\mathbf{k})=\begin{pmatrix}H_{\mathrm{U}}&&H_{\mathrm{Te}-\mathrm{U}}\\ H_{\mathrm{Te}-\mathrm{U}}^{\dagger}&&H_{\mathrm{Te}}\end{pmatrix}.$ (4)
@@
$\begin{pmatrix}&&H_{\mathrm{Te}-\mathrm{U}}\\ H_{\mathrm{Te}-\mathrm{U}}^{\dagger}&&\end{pmatrix}\mapsto\begin{pmatrix}&&\rho_{1}\;H_{\mathrm{Te}-\mathrm{U}}\\ H_{\mathrm{Te}-\mathrm{U}}^{\dagger}\;\rho_{1}&&\end{pmatrix},$ (5)
from which we read off that Eq. (3) is the only constant-in-momentum hybridization term that respects $\mathcal{M}_{z}$. Similarly, this term remains invariant under $\mathcal{M}_{y}$ symmetry, under which the hybridization transforms as $H_{\mathrm{Te-U}}\mapsto H_{\mathrm{Te-U}}\rho_{1}$.
[…]
@@
The parameters used for the normal state tight-binding band structure are displayed in Tab. 2. The minor changes of the parameters compared to Ref. [19] are motivated by recent low-energy tunneling measurements [38], and different from Ref. [19] we write the hybridization term in sublattice space. The main effect of these modified parameters is to produce slightly more dispersive Fermi surface sheets with dominant U content along $k_{z}$.
@@ II.2 Model for the superconducting state
with the $d$-vector describing generic triplet superconductivity; each component is a matrix in sublattice space $\rho_{i}$ and the space of U and Te as in Eq. (4). The symmetry of the system determines the possible pairing symmetry of the superconducting order parameter. As a result, the system will most likely condense into a superconducting phase characterized by an order parameter that transforms as one of the irreps of the symmetry group. Based on the current experimental findings on the superconducting state of UTe2, i.e. restricting to triplet superconductivity, the $d$-vector should transform as one of the odd-parity irreps of the $D_{2h}$ point group. In the limit of strong SOC, the four irreps exhibit the components as given in Table 3.
@@
Table 3: Possible forms of the $d$-vector for each symmetry-allowed odd-parity irrep of the $D_{2h}$ point group, assuming strong SOC (i.e. pseudospin and momentum transform together). Here, $X,Y,Z$ denote any function odd in $k_{x},k_{y},k_{z}$, respectively [46].
U sector Te sector U-Te sector $N_{\mathrm{terms}}$
$X$ $f_{x}\rho_{1/0}$, $f_{xz}\rho_{2}$ $f_{x}\rho_{1/0}$, $f_{xy}\rho_{2}$ $f_{x}(\rho_{0}+\rho_{1})$, $f_{y}(\rho_{0}-\rho_{1})$ 14
$Y$ $f_{y}\rho_{1/0}$, $f_{yz}\rho_{2}$ $f_{y}\rho_{1/0}$, $f_{1}\rho_{2}$ $f_{y}(\rho_{0}+\rho_{1})$, $f_{x}(\rho_{0}-\rho_{1})$ 14
$Z$ $f_{z}\rho_{1/0}$, $f_{1}\rho_{2}$ $f_{z}\rho_{1/0}$, $f_{zy}\rho_{2}$ $f_{z}(\rho_{0}+\rho_{1})$, $f_{xyz}(\rho_{0}-\rho_{1})$ 14
$XYZ$ $f_{xyz}\rho_{1/0}$, $f_{xy}\rho_{2}$ $f_{xyz}\rho_{1/0}$, $f_{xz}\rho_{2}$ $f_{xyz}(\rho_{0}+\rho_{1})$, $f_{z}(\rho_{0}-\rho_{1})$ 9
Table 4: Symmetry-allowed superconducting terms in sublattice space that transform according to the leftmost column. Here, $X$ is shorthand for a momentum-dependent function that transforms like $k_{x}$ under all point group symmetries of D2h and $f_{i}$ ($f_{xyz}$) are lattice harmonics that transform as $k_{i}$ ($k_{x}k_{y}k_{z}$).
@@
In the presence of finite hybridization, $\delta\neq 0$, the band transformation becomes more complicated. However, by inspecting the representations of the generators of the $D_{2h}$ point group in Tab. 1, one can verify that the superconducting terms written in Tab. 4 still transform correctly under all symmetry transformations. Finally, this also allows for the extension to superconducting terms in the hybridized U-Te sector written in the last column of Tab. 4, using the fact that $\rho_{0}+\rho_{1}$ is even while $\rho_{0}-\rho_{1}$ is odd under $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$ symmetries.
@@ II.3 Accessing the surface states
In the following, we analyze both the bulk and surface electronic properties of the superconducting phases of UTe2. In order to capture the surface states, we assume periodic boundary conditions in the two dimensions orthogonal to the surface plane, such that the parallel momentum $\mathbf{k}^{\parallel}$ remains a good quantum number. In the third direction (normal to the surface), the open boundary conditions break the translation invariance such that the full Hamiltonian can be written as
$H(\mathbf{k}^{\parallel})=\begin{pmatrix}H_{0}&V&0&\ldots&\\ V^{\dagger}&H_{0}&V&0&\ldots\\ 0&V^{\dagger}&H_{0}&\ddots&\\ \vdots&0&\ddots&\ddots&V\\ &\vdots&&V^{\dagger}&H_{0}\end{pmatrix}.$ (13)
The retarded Green’s function is given by
@@
where the highlighted blocks denote the inter-layer component on the two surfaces and in the bulk. Given the Hamiltonian form in Eq. (13), one can use the numerical method described in Ref. [47] to efficiently obtain $(G_{s},\tilde{G_{s}},G_{b})$. We refer to Appendix B for further details on how to extract these Green’s functions. From the surface and bulk Green’s functions, the observables of interest can be computed. For example, the electronic spectral density within a specific sector can be obtained by
@@
Finally, to properly describe the surface states on the experimentally relevant (0-11) cleave plane, we need to define a coordinate system that respects the symmetries of that specific surface. In particular, this requires introducing a new set of lattice vectors such that two of them leave the (0-11) surface invariant in addition to the crystal lattice. From this we define a surface parallel momentum $\mathbf{k}^{\parallel}=k_{x}\vec{m}_{x}+k_{c^{*}}\vec{m}_{c^{*}}$ in terms of $\vec{m}_{x}=\hat{e}_{x}/a$ and $\vec{m}_{c^{*}}=\hat{e}_{y}/b+\hat{e}_{z}/c$, where $a$, $b$ and $c$ are the lattice constants and $\hat{e}_{i}$, $i=x,y,z$, are the cartesian unit vectors. Due to the diagonal nature of the surface, the reciprocal lattice vectors are not proportional to the unit vectors in this coordinate system. Instead, they inherit the body-centered orthorhombic periodicity of the bulk unit cell such that the reciprocal lattice vectors have the forms $(2\pi,0)$ and $(\pi,2\pi)$, see Appendix A for more details. In Fig. 3 we show the resulting surface Brillouin zone projected onto the (0-11) surface and the reciprocal lattice vectors showing the periodicity of the BZ. We also plot the uranium part of the Fermi surface while leaving the tellurium sector out because it gives an almost constant background weight as a result of the flat dispersion of the bands.
@@ III.1 Bulk superconducting gap structure
To obtain the symmetry-allowed set of BdG Hamiltonians in each superconducting phase, we include all combinations with spin components according to Tab. (3) and momentum symmetries according to Tab. (4). We include momentum functions up to, and including, next-nearest neighbor coupling. Specifically, we use $f_{xy}(\mathbf{k})=\sin(k_{x}/2)\sin(k_{y}/2)\cos(k_{z}/2)$ and $f_{xyz}(\mathbf{k})=\sin(k_{x}/2)\sin(k_{y}/2)\sin(k_{z}/2)$ etc. For functions like $f_{x}(\mathbf{k})$ we have either $f_{x,1}(\mathbf{k})=\sin(k_{x}/2)\cos(k_{y}/2)\cos(k_{z}/2)$ corresponding to nearest-neighbor or $f_{x,2}(\mathbf{k})=\sin(k_{x})$, corresponding to next-nearest neighbor pairing. Including all symmetry-allowed terms results in 37 linearly independent terms for each phase with parameters that cannot be determined by symmetry only.
@@
Aside from the symmetry-imposed nodes discussed above, both the B2u and the B3u phases exhibit additional nodes that arise due to a combination of the symmetry of the order parameter and the body-centered orthorhombic periodicity of the BZ, as explained in detail below. The precise locations of these additional nodes are not imposed by symmetry. We do stress, however, that they are true nodes and not artifacts of including only the lowest-order pairing terms, i.e. including longer-range pairing terms will not gap them out.
[…]
Au $-1$ $-1$ $-1$ $-1$
B1u $1$ $1$ $-1$ $-1$
B2u $1$ $-1$ $1$ $-1$
B3u $-1$ $1$ $1$ $-1$
@@
To demonstrate this, consider the superconducting terms in the B2u phase in the $k_{z}=0$ plane. Here, the $\vec{d}$-vector only has one component $\vec{d}\sim(0,0,X)$ so it is the relevant plane to consider for the nodal structure of this order parameter. Since only one component is relevant at $k_{z}=0$, we consider the two types of momentum dependencies transforming as $X$ in the first row of Tab. (4). Assuming momentarily that the superconducting phase could avoid all but the symmetry-imposed nodes, this would require a dominant $f_{x,2}=\sin k_{x}$ pairing in the uranium sector since the nodal lines of $f_{x,2}$ have no overlap with the uranium Fermi surface, as seen in Fig. 4(a). A similar argument requires a dominant $f_{x,1}=\sin(k_{x}/2)\cos(k_{y}/2)\cos(k_{z}/2)$ pairing in the tellurium sector, as evident from Fig. 4(b). In the absence of any hybridization between the tellurium and uranium bands, this gap structure would result in a superconducting phase with only the symmetry-imposed nodes. However, at half of the band crossing points, the order parameter necessarily exhibits a different sign in the uranium and tellurium sectors, as seen by comparing Figs. 4(a,b) where four such points can be located. As a result of this sign change, any finite hybridization induces nodes at the avoided level crossing. In order to see this, consider a point in momentum space close to the band-crossing region. In that case the BdG Hamiltonian takes the simple form
$H=\begin{pmatrix}\xi&\delta&\Delta_{U}&0\\ \delta&\xi&0&\Delta_{Te}\\ \Delta_{U}&0&-\xi&-\delta\\ 0&\Delta_{Te}&-\delta&-\xi\\ \end{pmatrix},$ (16)
where $\xi_{U}=\xi_{Te}=\xi$ at the band crossing point, with $\xi_{U}$ and $\xi_{Te}$ denoting the U and Te bands in the absence of hybridization $\delta$. We are interested in regions of momentum space where the band crossing points exhibit opposite signs of the order parameter. Thus, let us assume simply $\Delta_{U}=-\Delta_{Te}=\Delta$. Transforming to a basis where the hybridized bands are diagonal, we obtain $\Delta\tau_{z}\rightarrow\Delta\tau_{x}$. Therefore, the hybridized bands necessarily feature a node near the band crossing point, which is in addition to the symmetry-imposed nodes along the high-symmetry $k_{y}$ axis. For a more general set of parameters, the nodes move to other positions of the Fermi surface as elaborated further below. We find that these nodes can only be removed by inter-band pairing terms which are unlikely to play a role for the tiny energy scales of the superconducting order relevant for UTe2.
@@
Figure 5: Overview of the positions of the symmetry-imposed and the additional nodes in the $k_{z}=0$ plane for B2u (a) and B3u (b) phases. In (a) relevant for B2u we display the gap function $d_{z,\mathrm{Te}},d_{z,\mathrm{U}}\sim(1-\alpha)f_{x,1}+\alpha f_{x,2}$ for four different values of $\alpha=0,0.2,0.5,1$. Panel (b) is the same as (a) except for B3u with the corresponding relevant gap function $d_{z,\mathrm{Te}},d_{z,\mathrm{U}}\sim(1-\alpha)f_{y,1}+\alpha f_{y,2}$. As seen, the symmetry-imposed nodes (marked by a square and triangle) remain fixed, whereas the additional nodes (marked by a star) move with $\alpha$ as indicated by the green arrows in both panels.

### v2

@@ I Introduction
The heavy-fermion compound UTe2 is currently a prime candidate for featuring an intrinsic spin-triplet superconducting phase in its ground state [3, 4, 5]. Indications for this include that the superconducting state survives in large magnetic fields exceeding the Pauli limiting field, and the weak temperature dependence of the Knight shift upon entering the superconducting phase [3, 6, 7, 8, 9, 10]. UTe2 is a body-centered orthorhombic material with D2h point group symmetry, allowing four symmetry-distinct spin-triplet order parameters in the case of strong spin-orbit coupling (SOC): Au, B1u, B2u, and B3u. Time-reversal symmetry can be broken below $T_{c}$ in the case of accidental near-degeneracies leading to superpositions of these states, as for example B1u + iB2u. At present, the experimental status of the nature of superconductivity in UTe2 remains controversial, with evidence in favor of all four allowed irreducible representations (irreps) and some also advocating for complex time-reversal symmetry-breaking non-unitary superpositions [11]. Thus, further experiments and theoretical modeling are needed to pinpoint the nature of the spin-triplet ordering of this material.
Several recent developments driven by improved sample quality appear to pave the way for significant progress in the understanding of superconductivity in UTe2 [12]. The superconducting state is not reached by multiple thermodynamic transitions and time-reversal symmetry remains preserved in the ground state condensate [13, 14]. It appears that the earlier signatures of time-reversal symmetry breaking below $T_{c}$ originated from extrinsic sources [15, 16, 17, 18]. Thus, the superconducting state most likely resides within a single of the above-mentioned irreps [19]. Recently, the normal state low-energy electronic structure and Fermi surface have been investigated in detail experimentally. Quantum oscillations are consistent with corrugated cylindrical Fermi sheets [20, 21]. Evidence for a closed Fermi pocket [22, 23, 24] is relatively weak, and hence we assume in this work the absence of closed Fermi pockets, which is important for the determination of the leading pairing instability and its associated topological properties [25, 26, 27, 28, 29]. Given the extensive evidence for nodal quasiparticles from e.g. thermal conductivity [30, 31], specific heat [32, 33] and penetration depth measurements [11], the leading candidates for the superconducting order parameter appear to be B2u and B3u.
@@ II Methodology
In this section, we first discuss the normal state band structure which is based on a slightly modified version of the density functional theory (DFT) calculations and subsequent four-band tight-binding model presented in Ref. 19. Second, we describe the derivation of the symmetry-allowed superconducting terms for both uranium (U) and tellurium (Te) sublattice sites. Finally, we outline the method used for extracting the surface Majorana bound states inside the superconducting gap.
@@ II.1 Model for the normal state
The starting point is a tight-binding model with few neighbor hoppings as used in Ref. 19 to describe the electronic structure and Fermi surface, consistent with recent conclusions from quantum oscillation measurements [20, 21]. The Hamiltonian of the U degrees of freedom can be expanded in Pauli matrices $\rho_{i}$ as
@@
$-\rho_{1}\;$ $[\Delta_{\mathrm{U}}+2t_{\mathrm{U}}^{\prime}\cos k_{x}+2t_{\mathrm{\rm ch,U}}^{\prime}\cos k_{y}$
@@
$\equiv\phantom{-}\rho_{0}\;$ $f_{\mathrm{U},0}(\mathbf{k})+\rho_{1}\;f_{\mathrm{U},1}(\mathbf{k})+\rho_{2}\;f_{\mathrm{U},2}(\mathbf{k}),\$(1)
@@
$\equiv\phantom{-}\rho_{0}\;$ $f_{\mathrm{Te},0}(\mathbf{k})+\rho_{1}\;f_{\mathrm{Te},1}(\mathbf{k})+\rho_{2}\;f_{\mathrm{Te},2}(\mathbf{k}).\$(2)
@@
$H_{N}(\mathbf{k})=\begin{pmatrix}H_{\mathrm{U}}&&H_{\mathrm{U}-\mathrm{Te}}\\ H_{\mathrm{U}-\mathrm{Te}}^{\dagger}&&H_{\mathrm{Te}}\end{pmatrix}$.\ (4)
@@
$\begin{pmatrix}&&H_{\mathrm{U}-\mathrm{Te}}\\ H_{\mathrm{U}-\mathrm{Te}}^{\dagger}&&\end{pmatrix}\mapsto\begin{pmatrix}&&\rho_{1}\;H_{\mathrm{U}-\mathrm{Te}}\\ H_{\mathrm{U}-\mathrm{Te}}^{\dagger}\;\rho_{1}&&\end{pmatrix}$, (5)
from which we read off that Eq. (3) is the only constant-in-momentum hybridization term that respects $\mathcal{M}_{z}$. Similarly, this term remains invariant under $\mathcal{M}_{y}$ symmetry, under which the hybridization transforms as $H_{\mathrm{U-Te}}\mapsto H_{\mathrm{U-Te}}\rho_{1}$.
$\mathcal{M}_{x}$ $\mathcal{M}_{y}$ $\mathcal{M}_{z}$ $\mathcal{I}$
@@
The parameters used for the normal state tight-binding band structure are displayed in Tab. 2. The minor changes of the parameters compared to Ref. 19 are motivated by recent low-energy tunneling measurements [38]. Distinct from Ref. 19 we write the hybridization term in sublattice space. The main effect of these modified parameters is to produce slightly more dispersive Fermi surface sheets with dominant U content along $k_{z}$.
@@ II.2 Model for the superconducting state
with the $\vec{d}$-vector describing generic triplet superconductivity; each component is a matrix in sublattice space $\rho_{i}$ and the space of U and Te as in Eq. (4). The symmetry of the system determines the possible pairing symmetry of the superconducting order parameter. As a result, the system will most likely condense into a superconducting phase characterized by an order parameter that transforms as one of the irreps of the symmetry group. Based on the current experimental findings on the superconducting state of UTe2, i.e. restricting to triplet superconductivity, the $\vec{d}$-vector should transform as one of the odd-parity irreps of the $D_{2h}$ point group. In the limit of strong SOC, the four irreps exhibit the components as given in Table 3.
@@
Table 3: Possible forms of the $\vec{d}$-vector for each symmetry-allowed odd-parity irrep of the $D_{2h}$ point group, assuming strong SOC (i.e. pseudospin and momentum transform together). Here, $X,Y,Z$ denote any function odd in $k_{x},k_{y},k_{z}$, respectively [46].
U sector Te sector U-Te sector
$X$ $f_{x}\rho_{1/0}$, $f_{xz}\rho_{2}$ $f_{x}\rho_{1/0}$, $f_{xy}\rho_{2}$ $f_{x}(\rho_{0}+\rho_{1})$, $f_{xyz}(\rho_{0}-\rho_{1})$
$Y$ $f_{y}\rho_{1/0}$, $f_{yz}\rho_{2}$ $f_{y}\rho_{1/0}$, $f_{1}\rho_{2}$ $f_{y}(\rho_{0}+\rho_{1})$, $f_{z}(\rho_{0}-\rho_{1})$
$Z$ $f_{z}\rho_{1/0}$, $f_{1}\rho_{2}$ $f_{z}\rho_{1/0}$, $f_{yz}\rho_{2}$ $f_{z}(\rho_{0}+\rho_{1})$, $f_{y}(\rho_{0}-\rho_{1})$
XYZXYZ $f_{xyz}\rho_{1/0}$, $f_{xy}\rho_{2}$ $f_{xyz}\rho_{1/0}$, $f_{xz}\rho_{2}$ $f_{xyz}(\rho_{0}+\rho_{1})$, $f_{x}(\rho_{0}-\rho_{1})$
Table 4: Symmetry-allowed superconducting terms in sublattice space that transform according to the leftmost column. Here, $X$ is shorthand for a momentum-dependent function that transforms like $k_{x}$ under all point group symmetries of D2h and $f_{i}$ ($f_{xyz}$) are lattice harmonics that transform as $k_{i}$ ($k_{x}k_{y}k_{z}$). $f_{1}$ transforms trivially under $D_{2h}$.
$i=1$ $i=2$
$f_{x,i}$ $\sin(k_{x}/2)\cos(k_{y}/2)\cos(k_{z}/2)$ $\sin k_{x}$
$f_{y,i}$ $\cos(k_{x}/2)\sin(k_{y}/2)\cos(k_{z}/2)$ $\sin k_{y}$
$f_{z,i}$ $\cos(k_{x}/2)\cos(k_{y}/2)\sin(k_{z}/2)$ $\sin k_{z}$
$f_{xy}$ $\sin(k_{x}/2)\sin(k_{y}/2)\cos(k_{z}/2)$ -
$f_{xz}$ $\sin(k_{x}/2)\cos(k_{y}/2)\sin(k_{z}/2)$ -
$f_{yz}$ $\cos(k_{x}/2)\sin(k_{y}/2)\sin(k_{z}/2)$ -
$f_{xyz}$ $\sin(k_{x}/2)\sin(k_{y}/2)\sin(k_{z}/2)$ -
Table 5: Lattice harmonics $f_{i}=f_{i}(\mathbf{k})$ corresponding to nearest and next-nearest neighbor pairing in the superconducting order parameter.
@@
In the presence of finite hybridization, $\delta\neq 0$, the band transformation becomes more complicated. However, by inspecting the representations of the generators of the $D_{2h}$ point group in Tab. 1, one can verify that the superconducting terms written in Tab. 5 still transform correctly under all symmetry transformations. Finally, this also allows for the extension to superconducting terms in the hybridized U-Te sector written in the last column of Tab. 5, using the fact that $\rho_{0}+\rho_{1}$ is even while $\rho_{0}-\rho_{1}$ is odd under $\mathcal{M}_{z}$ and $\mathcal{M}_{y}$ symmetries.
@@ II.3 Accessing the surface states
In the following, we analyze both the bulk and surface electronic properties of the superconducting phases of UTe2. In order to capture the surface states, we assume periodic boundary conditions in the two directions that lie in the surface plane, such that the parallel momentum $\mathbf{k}^{\parallel}$ remains a good quantum number. In the third direction (normal to the surface), the open boundary conditions break the translation invariance such that the full Hamiltonian can be written as
$H(\mathbf{k}^{\parallel})=\begin{pmatrix}H_{0}&V&0&\ldots&\\ V^{\dagger}&H_{0}&V&0&\ldots\\ 0&V^{\dagger}&H_{0}&\ddots&\\ \vdots&0&\ddots&\ddots&V\\ &\vdots&&V^{\dagger}&H_{0}\end{pmatrix}\;$, (13)
where $H_{0}$ describes the terms in the (BdG) Hamiltonian within a (suitably) chosen layer and $V$ is the coupling between nearest neighbor planes. The retarded Green’s function is given by
@@
where the highlighted blocks denote the inter-layer component on the two surfaces and in the bulk. Given the Hamiltonian form in Eq. (13), one can use the numerical method described in Ref. 47 to efficiently obtain $(G_{s},\tilde{G_{s}},G_{b})$. We refer to Appendix B for further details on how to extract these Green’s functions. Throughout this paper, we use $\eta=\Delta_{0}/300$ with $\Delta_{0}=10\>\mathrm{meV}$ where $\Delta_{0}$ is the overall prefactor of the superconducting terms. From the surface and bulk Green’s functions, the observables of interest can be computed. For example, the electronic spectral density within a specific sector can be obtained by
@@
Finally, to properly describe the surface states on the experimentally relevant (0-11) cleave plane, we need to define a coordinate system that respects the symmetries of that specific surface. In particular, this requires introducing a new set of lattice vectors such that two of them leave the (0-11) surface invariant in addition to the crystal lattice. From this we define a surface parallel momentum $\mathbf{k}^{\parallel}=k_{x}\vec{m}_{x}+k_{c^{*}}\vec{m}_{c^{*}}$ in terms of $\vec{m}_{x}=\hat{e}_{x}/a$ and $\vec{m}_{c^{*}}=\hat{e}_{y}/b+\hat{e}_{z}/c$, where $a$, $b$ and $c$ are the lattice constants and $\hat{e}_{i}$, $i=x,y,z$, are the Cartesian unit vectors. Due to the diagonal nature of the surface, the reciprocal lattice vectors are not proportional to the unit vectors in this coordinate system. Instead, they inherit the body-centered orthorhombic periodicity of the bulk unit cell such that the reciprocal lattice vectors have the forms $(2\pi,0)$ and $(\pi,2\pi)$, see Appendix A for more details. In Fig. 3 we show the resulting surface Brillouin zone projected onto the (0-11) surface and the reciprocal lattice vectors showing the periodicity of the BZ. We also plot the uranium part of the Fermi surface while leaving the tellurium sector out because it gives an almost constant background weight as a result of the particular orientation of the (0-11) plane and the rather flat dispersion of the Te bands.
@@ III.1 Bulk superconducting gap structure
To obtain the symmetry-allowed terms in the BdG Hamiltonian describing each of the superconducting phases, we include all combinations with spin components according to Tab. 3 and momentum symmetries according to Tab. 5. We include momentum functions up to, and including, next-nearest neighbor coupling as seen in Tab. 5.
@@
Aside from the symmetry-imposed nodes discussed above, both the B2u and the B3u phases exhibit additional nodes that arise due to a combination of the symmetry of the order parameter and the body-centered orthorhombic periodicity of the BZ, as explained in detail below. The precise locations of these additional nodes are not imposed by symmetry. We do stress, however, that they are true nodes and not artifacts of including only the lowest-order pairing terms, i.e. including longer-range pairing will not gap them out.
Irrep/$g$ $\mathcal{M}_{x}$ $\mathcal{M}_{y}$ $\mathcal{M}_{z}$ $\mathcal{I}$
Au 1-1 1-1 1-1 1-1
B1u 11 11 1-1 1-1
B2u 11 1-1 11 1-1
B3u 1-1 11 11 1-1
@@
To demonstrate this, consider the superconducting terms in the B2u phase in the $k_{z}=0$ plane. Here, the $\vec{d}$-vector only has one component $\vec{d}\sim(0,0,X)$ so it is the relevant plane to consider for the nodal structure of this order parameter. Since only one component is relevant at $k_{z}=0$, we consider the two types of momentum dependencies transforming as $X$ in the first row of Tab. 5. Assuming momentarily that the superconducting phase could avoid all but the symmetry-imposed nodes, this would require a dominant $f_{x,2}=\sin k_{x}$ pairing in the uranium sector since the nodal lines of $f_{x,2}$ have no overlap with the uranium Fermi surface, as seen in Fig. 4(a). A similar argument requires a dominant $f_{x,1}=\sin(k_{x}/2)\cos(k_{y}/2)\cos(k_{z}/2)$ pairing in the tellurium sector, as evident from Fig. 4(b). In the absence of any hybridization between the tellurium and uranium bands, this gap structure would result in a superconducting phase with only the symmetry-imposed nodes. However, at half of the band crossing points, the order parameter necessarily exhibits a different sign in the uranium and tellurium sectors, as seen by comparing Figs. 4(a,b) where four such points can be located. As a result of this sign change, any finite hybridization induces nodes at the avoided level crossing. In order to see this, consider a point in momentum space close to the band-crossing region. In that case the BdG Hamiltonian takes the simple form
$H=\begin{pmatrix}\xi&\delta&\Delta_{\rm U}&0\\ \delta&\xi&0&\Delta_{\rm Te}\\ \Delta_{\rm U}&0&-\xi&-\delta\\ 0&\Delta_{\rm Te}&-\delta&-\xi\\ \end{pmatrix}$, (16)
where $\xi_{\rm U}=\xi_{\rm Te}=\xi$ at the band crossing point, with $\xi_{\rm U}$ and $\xi_{\rm Te}$ denoting the U and Te bands in the absence of hybridization $\delta$. We are interested in regions of momentum space where the band crossing points exhibit opposite signs of the order parameter. Thus, let us assume simply $\Delta_{\rm U}=-\Delta_{\rm Te}=\Delta$. Transforming to a basis where the hybridized bands are diagonal, we obtain $\Delta\tau_{z}\rightarrow\Delta\tau_{x}$. Therefore, the hybridized bands necessarily feature a node near the band crossing point, which is in addition to the symmetry-imposed nodes along the high-symmetry $k_{y}$ axis. For a more general set of parameters, the nodes move to other positions of the Fermi surface as elaborated further below. We find that these nodes can only be removed by inter-band pairing terms which are unlikely to play a role for the tiny energy scales of the superconducting order relevant for UTe2.
@@
Figure 5: Overview of the positions of the symmetry-imposed and the additional nodes in the $k_{z}=0$ plane for B2u (a) and B3u (b) phases. In (a) relevant for B2u, we display the gap function $d_{z,\mathrm{Te}},d_{z,\mathrm{U}}\sim(1-\alpha)f_{x,1}+\alpha f_{x,2}$ for four different values of $\alpha=0,0.2,0.5,1$. Panel (b) is the same as (a) except for B3u with the corresponding relevant gap function $d_{z,\mathrm{Te}},d_{z,\mathrm{U}}\sim(1-\alpha)f_{y,1}+\alpha f_{y,2}$. As seen, the symmetry-imposed nodes (marked by a square and triangle) remain fixed, whereas the additional nodes (marked by a star) move with $\alpha$ as indicated by the green arrows in both panels.